Ordered: Digoxin 250mcg. Available: 0.125/10mL. How much medication should the patient receive (mL)?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

            Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

Here

      Desired dose(D) = 250 mcg.

      Dose in hand(H) = 0.125 mg.

Convert mg to mcg for calculation

        1 mg = 1000 mcg

So

     Dose in hand(H) = 0.125 * 1000

                                = 125 mcg.

         Quantity(Q) = 10 ml.

As per above formula

    Volume(ml) of drug required = D/H * Q

                                                  = 250/125 * 10

                                                  = 2 * 10

                                                  = 20 ml.